OpenStudy (anonymous):

d/dx [ln(x^3 + 1) + e^4x] How do I start this?

OpenStudy (math&ing001):

I'll give material that you'll need to DIY: \[\frac{ d \ln (A) }{ dx } = \frac{ \frac{ d A }{ dx } }{ A }\] \[\frac{ d e ^{A} }{ dx }=\frac{ d A }{ dx }*e ^{A}\]

OpenStudy (anonymous):

I don't understand how it is applied

OpenStudy (math&ing001):

\[\frac{ d \ln(x ^{3}+1) }{ dx }=\frac{ \frac{ d (x ^{3}+1) }{ dx } }{ x ^{3}+1 }=\frac{ 3x ^{2} }{ x ^{3}+1 }\] You do the same for d/dx(e^4x) and then you sum the results.